OpenStudy (anonymous):

Given the functions j(x) = x2 - 9 and k(x) = -x + 7, which operation results in a 3rd degree polynomial?

OpenStudy (anonymous):

@ganeshie8 @phi

OpenStudy (triciaal):

multiplication

OpenStudy (anonymous):

could you explain a little please? c:

OpenStudy (triciaal):

example : what do you have to do with 2 and 3 to get 6? multiply so 2 *3 = 6 what do you have to to with d and d to get d^2 multiply if you want to get x^3 and you have x^2 you need to multiply by x since x^3 = x^2 *x

OpenStudy (triciaal):

the question ask what you you have to do with (x^2-9) and (-x + 7) to get a term with x^3?